# NS-FPN: Improving Infrared Small Target Detection and Segmentation from Noise Suppression Perspective

### v1

@@ Abstract
Infrared small target detection and segmentation (IRSTDS) is a critical yet challenging task in defense and civilian applications, owing to the dim, shapeless appearance of targets and severe background clutter. Recent CNN-based methods have achieved promising target perception results, but they only focus on enhancing feature representation to offset the impact of noise, which results in the increased false alarms problem. In this paper, through analyzing the problem from the frequency domain, we pioneer in improving performance from noise suppression perspective and propose a novel noise-suppression feature pyramid network (NS-FPN), which integrates a low-frequency guided feature purification (LFP) module and a spiral-aware feature sampling (SFS) module into the original FPN structure. The LFP module suppresses the noise features by purifying high-frequency components to achieve feature enhancement devoid of noise interference, while the SFS module further adopts spiral sampling to fuse target-relevant features in feature fusion process. Our NS-FPN is designed to be lightweight yet effective and can be easily plugged into existing IRSTDS frameworks. Extensive experiments on the public IRSTDS datasets demonstrate that our method significantly reduces false alarms and achieves superior performance on IRSTDS tasks.
@@ 1 Introduction
Infrared small target detection and segmentation (IRSTDS) plays a critical role in various defense and civilian applications, including bird warning systems (Dai et al. 2021a), sea rescue operations (Yuan, Wang, and Wei 2022; Yuan and Wei 2024), and aerial surveillance (Zhang et al. 2024a; Yuan et al. 2024a). Due to the long imaging distance and lack of sufficient texture or structural information, infrared small targets (IRST) often appear as dim, shapeless spots, with low signal-to-noise ratios (SNR) and signal-to-clutter ratios (SCR) (Dai et al. 2023; Yuan et al. 2024b). These characteristics severely hinder accurate detection, particularly in dynamic environments where targets are buried in heavy background clutter and exhibit weak thermal signatures. Therefore, developing robust and efficient IRSTDS methods that can accurately localize small targets under diverse and realistic infrared scenarios remains a challenging problem.
Figure 1: Performance comparison of different methods on IRSTD-1K and NUAA-SIRST datasets. The outer regions represent superior performance.
Recent years have witnessed the research focus of IRSTDS shifting to CNN-based methods (Dai et al. 2021b, 2023; Yuan et al. 2024c), which concentrate on designing feature fusion structures to integrate high-level features with low-level details. For instance, Zhang et al. propose AGPCNet (Zhang et al. 2023) to perceive pixel correlations at specific scales for better feature representation. Concurrently, Dai et al. introduce OSCAR (Dai et al. 2023), a one-stage cascade refinement network, while Li et al. propose YOLOSR-IST (Li and Shen 2023) to introduce high-resolution feature maps into feature fusion. Beyond these detection-based methods, DNA-Net (Li et al. 2022) is explored to mitigate deep information loss caused by pooling layers, while Liu et al. (Liu et al. 2024) design a simple multi-scale head for the plain U-Net (MSHNet) to localize targets more precisely. Recently, Yuan et al. develop the SCTransNet (Yuan et al. 2024c) to achieve the encoding of multi-scale features while effectively addressing the challenge of modeling long-range dependencies in feature representation. Although these methods have achieved satisfactory performance improvements by designing complex network structures, they primarily focus on enhancing the feature representation while neglecting the false alarms introduced by the enhancement process. As shown in Figure 1, while the above methods exhibit satisfactory target localization performance in terms of $\mathrm{IoU}$ and $\mathrm{P_{d}}$, they concurrently exhibit high false alarm ($\mathrm{F_{a}}$) rates.
Figure 2: We employ the discrete haar wavelet transform (DWT) to decompose the original image into low and high frequency components and evaluate the performance of DNANet (Li et al. 2022) and MSHNet (Liu et al. 2024). We find that there is a crossover between the red and blue lines.
@@
High-frequency components are crucial for IRSTDS tasks while also lead to an increase in false alarm rates. Compared with the low-frequency components, the $\mathrm{P_{d}}$ and $\mathrm{IoU}$ metrics of both methods show excellent performance on the original images, which indicates the image details (high-frequency components) lost in the low-frequency components are crucial for improving target localization. However, these high-frequency components also contain noise interference, leading to an increase in false alarm rates.
@@
Inspired by the above, we pioneer in suppressing the noise in feature fusion to ensure improving target localization performance while reducing false alarm rate. Therefore, we firstly propose a low-frequency guided feature purification module (LFP) to suppress the noise features by purifying high-frequency components. Specifically, LFP begins with a 2D discrete wavelet transform (DWT) to decompose features into the frequency domain, then utilizes low-frequency features to predict the response at potential target locations. This prediction serves as a weighted map to refine high-frequency features, followed by gated gaussian filtering to further suppress less confident features. Finally, inverse DWT transformation yields noise-suppressed features while preserving high-frequency enhancement. Furthermore, to avoid interference from surrounding background noise, we design a spiral-aware feature sampling module (SFS) that performs spiral sampling in feature fusion process. Specifically, SFS employs dynamic sampling based on the intensity distribution characteristics (spiral shape) of IRST and calculates similarity to acquire the target-relevant features, further mitigating the impact of noise interference.
Different from previous methods that focus on designing complex network structures, our work aims to propose a lightweight yet effective way to improve performance. Thus, we integrate the LFP and SFS modules into the feature pyramid network (FPN) to construct a noise-suppression FPN, named NS-FPN. In this way, the LFP module can be used to replace the $1\times 1$ convolutions in FPN to achieve feature enhancement devoid of noise interference, while the SFS module substitutes the upsampling operations to achieve structured sampling and fusion of target-relevant features. These efficient designs allow our NS-FPN to be easily plugged into existing IRSTDS frameworks. Figure 1 shows that our NS-FPN can effectively reduce false alarms and achieve superior target localization performance. In summary, this paper makes the following contributions:
@@
➋ We propose a novel feature pyramid network (NS-FPN) that suppresses noise interference while enhancing target features through a low-frequency guided feature purification module and a spiral-aware feature sampling module.
@@ 2.1 IRST Detection and Segmentation Networks
With the advancement of deep learning techniques, CNN-based methods have emerged as the dominant paradigm, enabling automatic multi-layered feature learning for IRSTDS. For CNN-based detection, Dai et al. (Dai et al. 2023) proposed a one-stage cascade refinement network (OSCAR) that uses the results of high-level heads to modulate the predictions of low-level heads, completing the detection of a target from coarse to fine, and Yang et al. (Yang et al. 2024) introduced a dynamic head mechanism to adaptively adjust feature responses to different spaces and channels, thus improving the ability of network to focus on tiny targets. As for CNN-based segmentation, ISNet (Zhang et al. 2022) improved the accuracy of target shape prediction by introducing an edge reconstruction mechanism, while DNANet (Li et al. 2022) maintained the information of small targets in deep layers through dense interactions between features from the same and different layers. Recently, Liu et al. (Liu et al. 2024) used a novel scale and location sensitive (SLS) loss function to correct multi-scale prediction results, which can further improve the detection performance of existing methods. Concurrently, IRPruneDet (Zhang et al. 2024b) proposed a wavelet structure-regularized soft channel pruning strategy to improve the network inference efficiency.
@@
Figure 3: The overall structure of our NS-FPN and its conponents. Note that, convolutions are omitted for better visualization and $Y_{4}$ does not contains SFS module.
@@ 2.2 Feature Pyramid Network
Feature Pyramid Network (FPN) (Lin et al. 2017) emerged as a fundamental structure for general object detection, which can integrate high-level features with low-level details through a top-down connection. However, its linear summation may not be the most effective method for feature fusion. Based on it, path aggregation network (PANet) (Liu et al. 2018) improved the FPN by introducing a bottom-up path enhancement, which facilitates the reuse of low-level features and shortens information flow paths. Besides, bidirectional feature pyramid network (BiFPN) (Tan, Pang, and Le 2020) removed single-input nodes from PANet and introduced an adaptive weighted feature fusion mechanism to improve detection performance. In contrast to above FPNs that rely on manually crafted designs, NAS-FPN (Ghiasi, Lin, and Le 2019) leveraged neural architecture search to automatically explore the topological space of feature fusion networks and discovered optimal connection strategies.
In our work, we integrate LFP and SFS modules into the FPN structure to actively suppresses noise in feature fusion process, which is a specific design for the IRSTDS task.
@@ 3 Method
In this section, we introduce our noise-suppression feature pyramid network (NS-FPN) and its components. As illustrated in Figure 3, NS-FPN adopts a similar architectural design to conventional FPN, extracting hierarchical feature maps from the backbone network and employing $1\times 1$ convolutional operations to reduce their channel dimensions to 64. These channel-reduced features, denoted as $X_{1}$, $X_{2}$, $X_{3}$, $X_{4}$}, correspond to feature strides of {2, 4, 8, 16} pixels with respect to the original input resolution. The multi-scale feature pyramid $Y_{1}$, $Y_{2}$, $Y_{3}$, $Y_{4}$} is subsequently constructed by the top-down pathway within NS-FPN. Each lateral connection in NS-FPN integrate dual specialized modules for noise suppression: low-frequency guided feature purification (LFP) module and spiral-aware feature sampling (SFS) module. LFP suppresses the noise features by purifying high-frequency components in $X_{i}$ while SFS takes $X^{\prime}_{i}$, $Y_{i+1}$} as input, dynamically samples features in a spiral shape, and calculates similarity to further eliminate the influence of noise interference. Finally, the enhanced features $Y_{1}$, $Y_{2}$, $Y_{3}$, $Y_{4}$} obtained by feature addition are utilized for the subsequent IRSTDS tasks. Note that, all laterals of NS-FPN contain the LFP module, while only $Y_{1}$, $Y_{2}$, $Y_{3}$} layers contain the SFS module.
@@ 3.1 Low-frequency Guided Feature Purification
Figure 4: Detailed structure of the low-frequency guided feature purification (LFP) module.
Figure 5: Visualization of DAT and SFS sampling process.
@@ 3.2 Spiral-aware Feature Sampling
Specifically, for each LFP-purified feature map $X_{i}^{\prime}\in\mathbb{R}^{B\times C\times H_{i}\times W_{i}}$, we aim to inject higher-level semantic features $Y_{i+1}\in\mathbb{R}^{B\times C\times H_{i+1}\times W_{i+1}}$ from the deeper layer of NS-FPN. A straightforward way is to use DAT (Zhu et al. 2020) to randomly sample the $Y_{i+1}$ features to acquire scale-consistent features. However, since IRST is dim and small, random sampling cannot effectively capture the difference between the surrounding area and the target. Thus, we introduce an initialized pattern to constrain the sampling process. We first generate a uniform grid of reference points $p\in\mathbb{R}^{H_{G}\times W_{G}\times 2}$ for $Y_{i+1}$, where $H_{G}\times W_{G}$ controls the spatial sparsity of the sampling. Then, the offsets $\Delta p$ are generated by combining a basic spiral distribution $s$ and a group of learnable biases $\epsilon$:
@@
where $Attn(\cdot)$ is the cross attention (Carion et al. 2020) and $LN(\cdot)$ is LayerNorm (Ba, Kiros, and Hinton 2016). Finally, the output features $Y_{i}$ are obtained by modulating $X_{i}^{\prime}$ with $F_{s}$ through the residual addition:
@@ 3.3 Why the SFS Module Works?
Since using upsampling and pixel-by-pixel addition in the original FPN lacks spatial perception ability around IRST, a straightforward way is to adopt sparse attention mechanism (deformable attention (Zhu et al. 2020)) to enhance spatial perception while ensuring computational efficiency. However, as shown in Table 1, directly using DAT cannot effectively improve performance but increases computational complexity. The main reason is that IRST are typically small, occupy compact and shape-consistent regions, which makes the random sampling in sparse attention not applicable. Therefore, we propose the SFS module to effectively alleviate the above limitations from two aspects:
Spiral spatial perception. Since the intensity of IRST has the gaussian distribution characteristics, we explicitly restrict the sampling locations to satisfy the spiral distribution, as illustrated in Figure 5. This structured-design ensures the network can perceive fine-grained features around the IRST, which yields clearer target contours while cooperating with LFP to further suppress background noise interference and improve segmentation performance as shown in Table 1.
@@ 4.1 Datasets and Evaluation Metrics
Datasets. To evaluate the effectiveness of NS-FPN, we choose IRSTD-1k (Zhang et al. 2022) and NUAA-SIRST (Dai et al. 2021b) as our experimental datasets. IRSTD-1k includes 1000 real infrared images of $\times512$ in size, and NUAA-SIRST contains 427 infrared images of various sizes. For each dataset, 80% of the infrared images are used as training set and the remaining 20% are used as testing set.
Evaluation Metrics for Target Segmentation. For target segmentation tasks, Intersection over Union ($IoU$) is used as the pixel-level metric to evaluate shape description capability, and the probability of detection ($P_{d}$) and the false alarms ($F_{a}$) rate to evaluate localization performance.
Evaluation Metrics for Target Detection. For target detection tasks, mean Average Precision ($mAP$) is utilized to evaluate target detection performance based on classification accuracy and localization precision. Specifically, $mAP_{50}$ and $mAP_{75}$ denote the $mAP$ at a fixed $IoU$ threshold of 0.50 and 0.75 respectively, while $mAP$ is the average of $mAP$ values calculated over a range of $IoU$ thresholds from 0.50 to 0.95 with a step of 0.05.

### v2

# Seeing Through the Noise: Improving Infrared Small Target Detection and Segmentation from Noise Suppression Perspective

@@ Abstract
Infrared small target detection and segmentation (IRSTDS) is a critical yet challenging task in defense and civilian applications, owing to the dim, shapeless appearance of targets and severe background clutter. Recent CNN-based methods have achieved promising target perception results, but they only focus on enhancing feature representation to offset the impact of noise, which results in the increased false alarm problem. In this paper, through analyzing the problem from the frequency domain, we pioneer in improving performance from noise suppression perspective and propose a novel noise-suppression feature pyramid network (NS-FPN), which integrates a low-frequency guided feature purification (LFP) module and a spiral-aware feature sampling (SFS) module into the original FPN structure. The LFP module suppresses the noise features by purifying high-frequency components to achieve feature enhancement devoid of noise interference, while the SFS module further adopts spiral sampling to fuse target-relevant features in feature fusion process. Our NS-FPN is designed to be lightweight yet effective and can be easily plugged into existing IRSTDS frameworks. Extensive experiments on the IRSTD-1k and NUAA-SIRST datasets demonstrate that our method significantly reduces false alarms and achieves superior performance on IRSTDS task. The codes are available at https://github.com/mengduann/NS-FPN.
@@ 1 Introduction
Infrared small target detection and segmentation (IRSTDS) plays a critical role in various defense and civilian applications, including bird warning systems [5], sea rescue operations [28, 27], and aerial surveillance [32, 29]. Due to the long imaging distance and lack of sufficient texture or structural information, infrared small targets (IRST) often appear as dim, shapeless spots, with extremely low signal-to-noise ratios (SNR) and signal-to-clutter ratios (SCR) [7, 30]. These characteristics severely hinder accurate detection, particularly in dynamic environments where targets are buried in heavy background clutter and exhibit weak thermal signatures. Therefore, developing robust and efficient IRSTDS methods that can accurately localize small targets under diverse and realistic infrared scenarios remains a challenging problem.
Figure 1: Performance comparison of our method with other methods on the IRSTD-1k and NUAA-SIRST datasets. The outer regions represent superior performance.
Recent years have witnessed the focus of IRSTDS research shifting to CNN-based methods for detection [7, 36, 15] and segmentation [6, 13, 18, 31], which concentrate on designing feature fusion structures to integrate high-level features with low-level details. For instance, DNANet [13] is explored to mitigate deep information loss caused by pooling layers, and Liu et al. [18] design a simple multi-scale head for the plain U-Net (MSHNet) to localize targets more precisely. In addition, to capture the essential structural features of small targets, IRPruneDet [35] proposes a wavelet channel pruning method, which believes that high-frequency components can discriminate the importance of features for IRSTDS. Recently, IRSAM [34] improves the Perona-Malik diffusion equation [20] with a wavelet transform for image denoising and edge preservation, further improving detection performance. Although these methods have achieved satisfactory performance improvements by designing complex network structures, they primarily focus on enhancing the feature representation while neglecting the introduced false alarms, especially due to the emphasis on high-frequency components. As shown in Figure 1, while the above methods exhibit satisfactory target localization performance in terms of $\mathrm{IoU}$ and $\mathrm{P_{d}}$, they concurrently exhibit high false alarm ($\mathrm{F_{a}}$) rates in IRSTDS task.
Figure 2: The discrete haar wavelet transform (DWT) is utilized to decompose the original image into low and high frequency components. There is a crossover between the red and blue lines.
@@
High-frequency components are crucial for IRSTDS tasks while also leading to an increase in false alarm rates. Compared to low-frequency components, the $\mathrm{P_{d}}$ and $\mathrm{IoU}$ metrics of both methods show excellent performance on the original images, which indicates the image details (high-frequency components) lost in low-frequency components are crucial for improving target localization. However, these high-frequency components also contain noise interference, leading to an increase in false alarm rates.
@@
Inspired by the above, we pioneer in suppressing the noise in feature fusion to ensure improving target localization performance while reducing false alarm rates. Therefore, we firstly propose a low-frequency guided feature purification module (LFP) to suppress the noise features by purifying high-frequency components. Specifically, LFP begins with a 2D discrete wavelet transform (DWT) to decompose features into the frequency domain, then utilizes low-frequency features to predict the response at potential target locations. This prediction serves as a weighted map to refine high-frequency features, followed by gated gaussian filtering to further suppress less confident features. Finally, inverse DWT transformation yields noise-suppressed features while preserving high-frequency enhancement. Furthermore, to avoid interference from surrounding background noise, we design a spiral-aware feature sampling module (SFS) that performs spiral sampling in feature fusion. Specifically, SFS employs dynamic sampling based on the intensity distribution characteristics (spiral shape) and calculates similarity to acquire the target-relevant features, further mitigating the impact of noise interference.
Different from previous methods that focus on designing complex network structures, our paper aims to propose a lightweight yet effective way to improve performance. Thus, we integrate the LFP and SFS modules into the feature pyramid network (FPN) to construct a noise-suppression FPN, named NS-FPN. In this way, the LFP module can be used to replace the $1\times 1$ convolutions in FPN to achieve feature enhancement devoid of noise interference, while the SFS module substitutes the upsampling operations to achieve structured sampling and fusion of target-relevant features. These efficient designs allow our NS-FPN to be easily plugged into existing IRSTDS frameworks. Figure 1 shows that our NS-FPN can effectively reduce false alarms and achieve superior target localization performance. In summary, our contributions in this paper are highlighted as follows:
@@
➋We propose a novel feature pyramid network (NS-FPN) that suppresses noise interference while enhancing target features using a low-frequency guided feature purification module and a spiral-aware feature sampling module.
@@ 2.1 IRST Detection and Segmentation Networks
With the advancement of deep learning techniques, CNN-based methods have emerged as the dominant paradigm, enabling automatic multi-layered feature learning for IRSTDS. For CNN-based detection, Dai et al. [7] proposed a one-stage cascade refinement network (OSCAR) that uses the results of high-level heads to modulate the predictions of low-level heads, completing the detection of a target from coarse to fine, and Yang et al. [25] introduced a dynamic head mechanism to adaptively adjust feature responses to different spaces and channels, thus improving the ability of network to focus on tiny targets. As for CNN-based segmentation, ISNet [33] improved the accuracy of target shape prediction by introducing an edge reconstruction mechanism, while DNANet [13] maintained the information of small targets in deep layers through dense interactions between features from the same and different layers. Recently, Liu et al. [18] used a novel scale and location sensitive (SLS) loss function to correct multi-scale prediction results, which can further improve the detection performance of existing methods. Concurrently, IRPruneDet [35] proposed a wavelet structure-regularized soft channel pruning strategy to improve the network inference efficiency.
@@
Figure 3: The overall structure of our NS-FPN and the details of its conponents. For each scale, the feature is first fed into the LFP module, which generates a target-relevant spatial attention map based on low-frequency components to guide two-stage feature purification of high-frequency features. Subsequently, the purified feature is fused with the upper-layer feature through the SFS module to achieve semantic complementation according with spatial distribution prior of infrared small targets. Note that, convolutions are omitted for better visualization and $Y_{4}$ does not contain SFS module.
@@ 2.2 Feature Pyramid Network
Feature Pyramid Network (FPN) [16] emerged as a fundamental structure for general object detection, which can integrate high-level features with low-level details through a top-down connection. However, its linear summation may not be the most effective method for feature fusion. Based on it, path aggregation network (PANet) [19] improved the FPN by introducing a bottom-up path enhancement, which facilitates the reuse of low-level features and shortens information flow paths. Besides, bidirectional feature pyramid network (BiFPN) [22] removed single-input nodes from PANet and introduced an adaptive weighted feature fusion mechanism to improve detection performance. In contrast to above FPNs that rely on manually crafted designs, NAS-FPN [11] leveraged neural architecture search to automatically explore the topological space of feature fusion networks and discovered optimal connection strategies. Recently, HS-FPN [21] was designed for tiny object detection in visible images. However, it over-relied on high-frequency features determined by a preset spectral range and neglected the importance of low-frequency features, making it ineffective for IRST task. Unlike above methods, our work fully utilizes low-frequency components to adaptively guide high-frequency components, thereby actively suppressing noise in feature fusion process, which is a specific design for the IRSTDS.
@@ 3 Method
In this section, we introduce our noise-suppression feature pyramid network (NS-FPN) and its components. As illustrated in Figure 3, NS-FPN adopts a similar architectural design to conventional FPN, extracting hierarchical feature maps from the backbone network and employing $1\times 1$ convolutional operations to reduce their channel dimensions to 64. These channel-reduced features, denoted as $X_{1}$, $X_{2}$, $X_{3}$, $X_{4}$}, correspond to feature strides of {2, 4, 8, 16} pixels with respect to the original input resolution. The multi-scale feature pyramid $Y_{1}$, $Y_{2}$, $Y_{3}$, $Y_{4}$} is subsequently constructed by the top-down pathway within NS-FPN. Each lateral connection in NS-FPN integrates dual specialized modules for noise suppression: low-frequency guided feature purification (LFP) module and spiral-aware feature sampling (SFS) module. LFP suppresses the noise features by purifying high-frequency components in $X_{i}$ while SFS takes $X^{\prime}_{i}$, $Y_{i+1}$} as input, dynamically samples features in a spiral shape, and calculates similarity to further eliminate the influence of noise interference. Finally, the enhanced features $Y_{1}$, $Y_{2}$, $Y_{3}$, $Y_{4}$} obtained by feature addition are utilized for the subsequent IRSTDS tasks. Note that, all laterals of NS-FPN contain the LFP module, while only $Y_{1}$, $Y_{2}$, $Y_{3}$} layers contain the SFS module.
@@ 3.2 Spiral-aware Feature Sampling
Specifically, for each LFP-purified feature map $X_{i}^{\prime}\in\mathbb{R}^{B\times C\times H_{i}\times W_{i}}$, we aim to inject higher-level semantic features $Y_{i+1}\in\mathbb{R}^{B\times C\times H_{i+1}\times W_{i+1}}$ from the deeper layer of NS-FPN. A straightforward way is to use DAT [37] to randomly sample the $Y_{i+1}$ features to acquire scale-consistent features. However, since IRST is dim and small, random sampling cannot effectively capture the difference between the surrounding area and the target. Thus, we introduce an initialized pattern to constrain the sampling process. We first generate a uniform grid of reference points $p\in\mathbb{R}^{H_{G}\times W_{G}\times 2}$ for $Y_{i+1}$, where $H_{G}\times W_{G}$ controls the spatial sparsity of the sampling. Then, the offsets $\Delta p$ are generated by combining a basic spiral distribution $s$ and a group of learnable biases $\epsilon$:
@@
where $Attn(\cdot)$ is the cross attention [3] and $LN(\cdot)$ is LayerNorm [1]. Finally, the output features $Y_{i}$ are obtained by modulating $X_{i}^{\prime}$ with $F_{s}$ through the residual addition:
@@ 3.3 Why the SFS Module Works?
Since using upsampling and pixel-by-pixel addition in the original FPN lacks spatial perception ability around IRST, a straightforward way is to adopt deformable attention (DAT) [37] to enhance spatial perception while ensuring computational efficiency. However, as shown in Table 1, directly using DAT cannot effectively improve performance but increases computational complexity. The main reason is that IRST are typically small, occupy compact and shape-consistent regions, which makes the random sampling in sparse attention not applicable. Therefore, we propose the spiral-aware feature sampling to effectively alleviate the above limitations from two aspects:
Spiral spatial perception. Since the intensity of IRST has the gaussian distribution characteristics, we explicitly restrict the sampling locations to satisfy the spiral distribution, as illustrated in Figure 4. This structured-design ensures the network can perceive fine-grained features around the IRST, which yields clearer target contours while cooperating with LFP to further suppress background noise interference as shown in Table 1.
@@ 4.1 Datasets and Evaluation Metrics
Datasets. To evaluate the effectiveness of NS-FPN, we choose IRSTD-1k [33] and NUAA-SIRST [6] as our experimental datasets. IRSTD-1k includes 1000 real infrared images of $\times512$ in size, and NUAA-SIRST contains 427 infrared images of various sizes. For each dataset, 80% of the infrared images are used as training set and the remaining 20% are used as testing set.
Evaluation Metrics for Segmentation. For segmentation tasks, Intersection over Union ($IoU$) is used as the pixel-level metric to evaluate shape description capability, and the probability of detection ($P_{d}$) and the false alarms ($F_{a}$) rate to evaluate localization performance.
Evaluation Metrics for Detection. For detection tasks, mean Average Precision ($mAP$) is utilized as the primary metric to evaluate detection performance based on classification accuracy and localization precision. Specifically, $mAP_{50}$ and $mAP_{75}$ denote the $mAP$ at a fixed $IoU$ threshold of 0.50 and 0.75 respectively, while $mAP$ is the average of $mAP$ values calculated over a range of $IoU$ thresholds from 0.50 to 0.95 with a step of 0.05.
4.2 Implementation Details
All experiments are conducted on NVIDIA GeForce RTX 4090 GPUs. We integrate NS-FPN into MSHNet [18] for segmentation and YOLOv8n-p2 for detection. The Adagrad optimizer [9] is used with an initial learning rate of 0.05. The models are trained for 500 epochs with a batch size of 16. For segmentation tasks, input images are resized to 256 and then randomly cropped to 224 during training. For detection tasks, input images are uniformly resized to 640.